### v1

# State Permutation Control in Non-Hermitian Multiqubit Systems with Suppressed Non-Adiabatic Transitions

Abstract
Non-Hermitian systems have been at the focus of intense research for over a decade, partly due to their nontrivial energy topology formed by intersecting Riemann manifolds with branch points known as exceptional points (EPs). This spectral property can be exploited, e.g., to achieve controlled state permutations that are necessary for implementing a wide class of classical and quantum information protocols. However, the imaginary spectra of typical non-Hermitian systems lead to instabilities and breakdown of adiabaticity, which impedes the practical use of EP-induced energy topologies in quantum information protocols that explicitly rely on symmetric state flips. On the other hand, techniques that help to suppress non-adiabatic transitions in non-Hermitian systems have also been developed, but these advances have so far been limited to single qubits. In this work, we address this long-standing problem by introducing a model of interacting qubits governed by an effective non-Hermitian Hamiltonian that hosts EPs and possesses a completely real energy spectrum. We demonstrate that such non-Hermitian Hamiltonians enable realization of permutation groups in the multi-qubit eigenspace. Our findings indicate that, contrary to previous beliefs, non-Hermiticity can be utilized to achieve controlled state permutations in time-modulated multiqubit systems, thus paving the way for the advancement and development of novel quantum information protocols.
I Introduction
The presence of Riemann topology with branch points, called exceptional points (EPs), in the energy spectrum of non-Hermitian systems has attracted considerable interest over the last few decades. The EPs, where non-Hermitian Hamiltonian (NHH) eigenvectors coincide, have been linked to a number of exciting and nontrivial effects and phenomena not found in Hermitian systems [1, 2, 3, 4].
From the early days of research on EPs, it has been suggested that the topology of the intersecting Riemann energy manifolds at EPs can be exploited for the implementation of a symmetric state switch [5, 6, 7, 8, 9]. In other words, by dynamically encircling EPs in the system parameter space, an initial eigenstate, belonging to one of the Riemann energy sheets, adiabatically transitions to higher or lower-lying Riemann energy eigenstates, thereby enabling a symmetric state transfer [10, 11, 12, 13, 14, 15, 16].
However, subsequent studies revealed that symmetric state switching via dynamical encirclement of EPs is generally impractical due to the breakdown of adiabaticity in non-Hermitian systems, driven by the presence of an imaginary spectrum [17, 18, 19, 20]. In particular, any initial eigenmode of the NHH evolves into the state with the minimal loss through a series of non-adiabatic transitions (NATs) during the system dynamics. This intriguing chiral mode behavior has also been demonstrated in experiments [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], and various approaches for state control have been explored in the context of such quantum dynamics [33, *Ibanez_2014, 35, 36, 37, 38, 39, 40]. Although some progress has been made in restoring symmetric state flips to enable a programmable multi-state switch [41], the proposed approach does not fully recover the symmetry of the state transfer, as it still depends on the occurrence of NATs in the system.
In a recent work in Ref. [42] some of us demonstrated that this seeming difficulty can be resolved in two-level systems by carefully choosing state trajectories in the system’s parameter space, allowing adiabaticity to be fully restored while dynamically winding around EPs. However, whereas the method can be potentially extended to linear optical multimode systems by combining the approach with findings in Ref. [41], it presents a significant challenge when applied to interacting multiqubit systems.
In this work, we address this long-standing problem by introducing an NHH that describes an interacting multiqubit system and exhibits a purely real spectrum with EPs defined on hyperboloid manifolds in the system parameter space. Remarkably, by dynamically encircling the EPs on such manifolds, it enables the implementation of a genuine and controllable symmetric multi-state switching; thus facilitating the generation of various permutation groups in the NHH eigenspace, a property that had been previously thought challenging or impossible to achieve.
The ability to generate and control permutation groups in the NHH eigenspace holds promise for new opportunities in quantum information and communication. Potential applications include the production of diverse entangled states [43], the implementation of optimal quantum gates [44], the development of quantum error-correcting codes [45, 46], performing holonomic quantum computations [47], and advancements in quantum sensing [48], among others.
As an illustrative example, we study here an NHH describing a three-qubit system and investigate the rich structure of eigenstate permutations that arise on dynamically orbiting its associated EPs. Our analysis reveals that the symmetric state switching protocol implemented here can generate a nontrivial non-Abelian permutation group comprising 576 elements. We also show that the proposed approach demonstrates robustness against external noise and perturbations, and it can be extended to any number of qubits. Our work therefore opens avenues for advancing and developing novel quantum information protocols, in which generating and steering permutations of quantum states plays a pivotal role.
II Model
We start by introducing a tight-binding non-Hermitian Hamiltonian $\hat{H}\neq\hat{H}^{\dagger}$ describing $N$ qubits represented as spins with $ZZ$ coupling and which are subject to a transverse, in general non-uniform, effective complex magnetic field
$\hat{H}=\sum\limits_{k}^{N}\vec{{\boldsymbol{\sigma}}}^{k}\vec{\mathbb{B}}_{||}^{k}+\sum\limits_{m<n}J_{mn}\hat{\sigma}_{z}^{m}\hat{\sigma}_{z}^{n}.$ (1)
Here, the spin vector $\vec{{\boldsymbol{\sigma}}}=[\hat{\sigma}_{x},\hat{\sigma}_{y},0]^{T}$, with the symbol $T$ denoting the transposed operation, $\hat{\sigma}_{x,y,z}$ are the Pauli matrices, and the magnetic field vector is written as
$\vec{\mathbb{B}}_{||}^{k}=\Big{[}\alpha_{k}\sin\phi,\alpha_{k}\cos\phi,0\Big{]}^{T},$ (2)
where
$\alpha_{k}=f_{k}\frac{y\sin\left({\rm Re}[\phi]\right)}{\sin\left({\rm Im}[\phi]\right)},\quad\phi=\arctan\left[(x+iy)^{-1}\right],$ (3)
with $f_{k},x,y\in{\mathbb{R}}$. For fixed $f_{k}$, the system parameter space is defined by a tuple $(x,y,J)$, with $J$ being the coupling matrix $J\equiv J_{mn}$.
The equations (2) and (3) determine the embedding map $(x,y)\to(B_{x}^{r},B_{x}^{i},B_{y}^{r},B_{y}^{i})$, where superscripts in $B$ denote real and imaginary parts of the magnetic field. The components of the local magnetic field, thus, describe a hyperboloid surface with a varied curvature $\alpha_{k}$ in ${\mathbb{R}}^{4}$ [42], namely
$(B_{x}^{r})^{2}+(B_{y}^{r})^{2}-(B_{x}^{i})^{2}-(B_{y}^{i})^{2}=\alpha_{k}^{2}.$ (4)
This parametrization of the magnetic field, together with the $ZZ$ qubit interactions, ensures that the NHH in Eq. (1) is pseudo-Hermitian and this symmetry is exact in the whole parameter space, meaning the spectrum of $\hat{H}$ is everywhere real. Indeed, the condition for the pseudo-Hermiticity of the NHH reads as [49]
$\eta\hat{H}\eta^{-1}=\hat{H}^{\dagger},$ (5)
where $\eta$ is some Hermitian operator. Here, the operator $\eta$ can be chosen as a metric operator of the NHH describing non-interacting qubits, i.e., when $J=\boldsymbol{0}$ (see also Appendix A).
The NHH in Eq. (1) can be implemented in various experimental platforms, ranging from solid-state systems to photonic setups, utilizing either the Naimark dilation method [50] or photon postselection techniques [51]. We elaborate on this in more detail in Sec. V.
III Eigenspectrum of the non-Hermitian Hamiltonian and ‘twisted’ exceptional points
Here we discuss the spectral properties of the NHH in Eq. (1), and the potential of the system with an odd number of qubits to realize a genuine permutation subgroup.
III.1 Non-interacting qubits
We begin our analysis of the spectral properties of the NHH in Eq. (1) with the simplest case, when $J\equiv\boldsymbol{0}$, and $f_{k}=1$, for $\forall k$. In that case, the spectrum of $N$ non-interacting qubits is analogous to that of an $N$-dimensional hypercube [52, 53], namely
$E_{k}\equiv-N\alpha,-(N-2)\alpha,\dots,(N-2)\alpha,N\alpha,$ (6)
which is generated by the spectrum of a single qubit $E^{\rm q}_{1,2}=\pm\alpha$. The latter we plot in Fig. 1. In what follows, we assume that the energy levels are always arranged in ascending order, i.e., $E_{1}\leq E_{2}\leq\dots\leq E_{2N}$.
Figure 1: Energy spectrum of a single qubit, described by the NHH in Eq. (1), as a function of the applied magnetic field $\vec{\mathbb{B}}_{||}^{i}(x,y)$, according to Eqs. (6) and (3). The purely real-valued spectrum consists of two intertwined Riemann sheets with an EP located at $(x=0,y=1)$. Moreover, at this EP the spectrum of the NHH diverges, according to Eq. (8) (see the main text for details).
The right eigenvector space of the non-interacting qubits’ NHH is spanned by tensor products of the eigenvectors of each individual qubit that read [52, 42]
$|\psi_{1}\rangle\equiv\hat{R}\begin{pmatrix}-\sin\dfrac{\phi}{2}\\ \cos\dfrac{\phi}{2}\end{pmatrix},\quad|\psi_{2}\rangle\equiv\hat{R}\begin{pmatrix}\cos\dfrac{\phi}{2}\\ \sin\dfrac{\phi}{2}\end{pmatrix},$ (7)
where $\hat{R}=\exp\left(-i\pi\sigma_{x}/4\right)$ is a rotation operator. Because the pseudo-Hermiticity of the parameterized NHH does not experience spontaneous symmetry breaking, the left eigenvectors are simply the complex conjugates of the right eigenvectors [42].
We now proceed with a more detailed examination of the spectral properties of a single qubit, as its properties are directly reflected in Eq. (6). As can be seen from Fig. 1, the qubit spectrum is characterized by the presence of a singular point at $(x=0,y=1)$. Indeed, according to Eq. (3) one has
$\lim\limits_{x\to 0,y\to 1}\alpha=\pm\infty,$ (8)
where the sign at infinity depends on the direction of the limit approach to the singularity (see also Fig. 1). Furthermore, one finds the limit for the inner product of the normalized eigenvectors at this point to be
$\lim\limits_{x\to 0,y\to 1}|\langle\psi_{2}|\psi_{1}\rangle|=1.$ (9)
Equation (9) implies that the qubit eigenvectors coalesce at this singular point. In other words, the singularity behaves like an EP, though it is not an EP in the usual sense that is encountered in systems with finite spectra. This seeming difficulty can be handled using mathematical tools borrowed from the projective space framework. Put differently, this singularity can be cast into an ordinary EP in some extended NHH, where the eigenvalues converge at the EP (see Appendix C for details).
Nevertheless, for our purposes here, we can safely ignore the infinite nature of the spectrum at such EPs, because we are solely interested in the encircling trajectories around the EPs, along which both the energy spectrum and the system controlling parameters (magnetic fields and interactions) remain finite. These EPs may therefore be termed as ’twisted EPs,’ due to the nontrivial wrapping of the spectrum at these points that leads to the divergent behavior of the NHH eigenvalues. For simplicity, however, we will continue to refer to such singularities as EPs.
When dynamically encircling the EP in Fig. 1 in the parameter space $(x,y)$, the eigenstates of a qubit can be interchanged adiabatically without invoking any non-adiabatic transitions [42]. This stems from the fact that the eigenvalues are everywhere real. Note that while a winding trajectory around the EP in the space $(x,y)$ can have the form of simple loops in a plane, the actual orbits in the vector space (${\mathbb{C}}^{2}$ or ${\mathbb{R}}^{4}$) of the magnetic field $\vec{\mathbb{B}}_{||}$ are the loops on a curved hyperboloid.
We note that the spectrum described above drastically differs from that studied in Ref. [42], where the qubit eigenvalues have a well-defined limit in Eq. (8). The key distinction lies in the fact that, unlike in Ref. [42], the Riemann energy sheets are well separated at the point $(x=0,y=0)$ where the NHH becomes Hermitian (see Fig. 1). This is a very useful property, as it can allow performing state transfer protocols over multiqubit states which are often generated and exploited in closed systems. The latter thus enables the combination and extension of existing quantum information protocols, originally devised for Hermitian systems, to non-Hermitian systems as well.
Figure 2: Eigenspectrum of the NHH $\hat{H}$ in Eq. (1) as a function of the magnetic field $\vec{\mathbb{B}}_{||}^{i}(x,y)$ for (a) two interacting qubits and (b) three interacting qubits. The system parameters are (a) $f_{2}=2f_{1}=2$, $J_{12}=1$; and (b) $f_{1,2,3}=1$, $J_{12}=0.5$, $J_{13}=1$; $J_{23}=1.5$. The spectrum consists of a stack of Riemann sheet pairs, whose number grows with the number of interacting qubits in the system. The set of EPs, constituting the exceptional line (EL), is located at the same point $(x=0,y=1)$ as in the case of a single non-interacting qubit in Fig. 1.
III.2 Interacting qubits
When the qubit interactions in Eq. (1) are switched on, the NHH spectrum takes the form of a stack of Riemann sheet pairs, as illustrated in Fig. 2, with the spacing between sheets determined by a function of the $ZZ$ coupling strength $J_{mn}$. Hence, for interacting qubits, the EP of a single qubit extends to an exceptional line (EL), as depicted in Fig. 2. This EL is orthogonal to the $(x,y)$ plane and intersects it at the point $(x=0,y=1)$.
To determine whether a non-Hermitian multiqubit system enables a multistate switch, one can examine a topological invariant, associated with EPs, known as chirality [13]. Namely, when the chirality is non-zero, the eigenstates can be swapped while encircling an EP. In contrast, when the chirality is zero, winding around the EP does not induce a state flip; instead, each eigenstate returns to itself after completing a full encircling cycle [13].
A non-zero chirality can be inferred from the behavior of the spectrum near the Riemann branch cuts where $\alpha=0$, which corresponds to the region $x=0,y\in[1,\infty)$ in the $(x,y)$ plane (see Figs. 1 and 2). Indeed, at the branch cut, the function $\alpha$ changes sign. This means that if a decomposition of $E_{k}$ in this region contains odd powers of $\alpha$, this will be an indication that the levels interchange when crossing a branch cut.
One can readily verify that, in general, for arbitrary values of magnetic fields $\vec{\mathbb{B}}_{||}^{i}$ and coupling strengths $J_{mn}$, the energy $E_{k}$, ($k=1,\dots,2N$), in the vicinity of each branch cut can be expanded as follows
$E_{k}=\begin{cases}\sum\limits_{j}u_{j}^{k}\alpha^{j},\quad&\text{when $N$ is odd},\\ \sum\limits_{j}u_{j}^{k}\alpha^{2j},\quad&\text{when $N$ is even},\end{cases}$ (10)
with $u_{j}^{k}\in{\mathbb{R}}$. From Eq. (10) one concludes that a non-zero chirality of the EPs and state transfer, associated with it, are always realizable in a system with an odd number of qubits.
We note that a non-zero chirality, and therefore the symmetric state transfer, for an even number of qubits can also be realized by making appropriate modifications to the NHH in Eq. (1). For example, in a two-qubit system, this can be achieved by applying a non-Hermitian synthetic magnetic field to one qubit, while the second qubit is subjected to a real-valued field.
IV Implementing a state permutation group in a non-hermitian time-modulated three-qubit system
Figure 3: Schematic representation of a three-qubit spin system with pairwise $ZZ$ interactions, subject to a transverse magnetic field (simplified here as a perpendicular vector to the plane). The system is governed by the effective NHH in Eq. (1). The magnetic field strength acting on each spin qubit generally varies, as indicated by the differently colored clouds surrounding the qubits. By appropriately time-modulating magnetic fields and qubit interactions, according to Eq. (IV.1), one can realize a nontrivial permutation group acting on the system’s NHH eigenspace.
We now proceed to describe the implementation of a nontrivial state permutation group in the eigenspace of a time-modulated three-qubit system as an illustrative example (see also Fig. 3).
IV.1 Description of the state transfer protocol
Figure 4: Dynamically encircling EPs in a three-qubit system governed by the NHH in Eq. (1) by time modulating the magnetic field $\vec{\mathbb{B}}_{||}^{i}(x,y)$, according to Eq. (IV.1), over a winding period $T$. (a) A schematic representation of the initial eigenstates at $t=0$, depicted as colored spheres, with each color corresponding to a certain Riemann energy manifold. (b),(c) The evolution of various initialized states at times $T/2<t<T$ and at $t=T$, respectively. The Riemann energy sheet configuration shown in panel (b) corresponds to $t=T/2$. (d) Fidelity $|\langle\psi_{k}|\psi(t)\rangle|^{2}$ between the evolving state $|\psi(t)\rangle$ and the initial eigenstate $|\psi_{k}\rangle$, $k=1,\dots,8$, of the NHH. System parameters are set as $f_{3}=2f_{1}=2f_{2}=2$, $J_{12}=J_{13}=J_{23}=1$. The time period is $T=1500$ [arb. units], the encircling radius $r=1$, and the initial phase $\phi_{0}=\pi$.
Figure 5: Dynamically spiraling EPs in a three-qubit system governed by the NHH in Eq. (1) by time modulating both the magnetic field $\vec{\mathbb{B}}_{||}^{i}(x,y)$ and qubit coupling $J_{13}$, according to Eq. (IV.1), over a winding period $T$. (a) A schematic representation of the initial eigenstates at $t=0$, depicted as colored spheres, with each color corresponding to a certain Riemann energy manifold. (b),(c) The evolution of various initialized states at times $T/2<t<T$ and at $t=T$, respectively. The Riemann energy sheet configuration shown in panel (b) corresponds to $t=T/2$, which is different from that in panel (a), and which is induced by the additional modulation of the qubit coupling $J_{13}(t)$. The latter enables to couple different Riemann energy sheets compared to that in Fig. 3 with static interactions. (d) Fidelity $|\langle\psi_{k}|\psi(t)\rangle|^{2}$ between the evolving state $|\psi(t)\rangle$ and the initial eigenstate $|\psi_{k}\rangle$, ($k=1,\dots,8$), of the given three-qubit NHH. The rest of the system parameters are the same as in Fig. 3. We note that the same result is obtained when modulating the qubit coupling $J_{23}$ in Eq. (IV.1).
In Fig. 2(b) we present a typical energy spectrum for the interacting three qubits, governed by the NHH in Eq. (1). As can be seen from the figure, the winding around EPs in the $(x,y)$ plane alone can only result in the state transfer between eigenstates belonging to individual Riemann pairs. In order to achieve the state flip between different Riemann pairs as well, one has to ensure that these sheets can be ‘transported’ along the $z$-axis in Fig. 2(b). The intended approach is reminiscent to the case studied in Ref. [41], where individual pairs of Riemann manifolds of a bosonic multimode spectrum were manipulated in a similar manner. Here, such control can be achieved by further modulating the interaction terms $J_{mn}$ in the NHH. Specifically, we require the system parameters to be time-modulated as follows:
$x(t)$ $=$ $r\sin(\omega t+\phi_{0}),$
$y(t)$ $=$ $1+r\cos(\omega t+\phi_{0}),$
$J_{mn}(t)$ $=$ $J_{mn}\cos^{2}({\omega t}/{2}),$ (11)
where $r$ is the circle radius in the $(x,y)$ plane, $\omega=2\pi/T$ is an angular frequency determined by a winding period $T$, and $\phi_{0}$ is the initial phase. Moreover, the qubit coupling strengths $J_{mn}$, with $m,n=1,2,3$, and $m<n$, can be modulated individually, i.e., not necessarily simultaneously. To track a state evolution in such a time-modulated system, we numerically solve the Schrödinger equation $i\partial_{t}|\psi(t)\rangle=H|\psi(t)\rangle$ with the Runge-Kutta method.
Alternatively, one could describe the trajectory in Eq. (IV.1) as dynamical spiraling around EPs rather than mere dynamical encirclement. This is because, in addition to orbiting a circular path in the 2D $(x,y)$ parameter space, the system is simultaneously driven along a perpendicular axis, which is defined by the couplings $J_{mn}$, as illustrated in Fig. 2.
IV.2 Dynamical encirclement and spiraling around exceptional points in a three-qubit system
To elucidate the state transfer characterized by the overlap of distinct Riemann pairs during the dynamical encirclement and spiraling of EPs, we give illustrative examples in Figs. 4 and 5, respectively.
When plotting Fig. 4, it is assumed that time modulation is only performed over the parameters $(x,y)$, leaving interactions unchanged $J_{12}=J_{13}=J_{23}=1$, i.e., we solely implement the dynamical encirclement of EPs.
We initialize the systems in one of the NHH eigenstates, as depicted by colored spheres in Fig. 4(a). By starting orbiting around EPs in the $(x,y)$ plane, at the time $t=T/2$ the evolving states cross corresponding energy branch cuts resulting in the state transfer, as shown in panels (b) and (c) in Fig. 4. We calculate the fidelity between the evolving state $\psi(t)$ and initial eigenstates of the NHH at $t=0$ in Fig. 4(d) when EPs encirclement is realized in the counterclockwise direction. We note that identical results are obtained when winding in the opposite, i.e., clockwise, direction, because the non-adiabatic jumps are eliminated in the system due to the real spectrum of the NHH. The resulting permutation can be symbolically written in a cyclic notation as $p=(1,5)(2,6)(3,4)(7,8)$, where the numbers in parentheses account for the cyclic permutations between corresponding eigenvectors. For instance, the notation $(1,5)$ means that the pair of eigenstates $\psi_{1}$ and $\psi_{5}$ are permuted with each other.
It is evident that with the fixed qubit couplings described above, the system is significantly limited in the number of realizable eigenstate flips. To overcome this restriction one may additionally modulate the coupling between, e.g., the first and third qubits, i.e., by modulating the parameter $J_{13}(t)$ in Eq. (IV.1).
We demonstrate the outcome of such a dynamical spiraling around EPs in Fig. 5. The set of possible initialized eigenstates in Fig. 5(a) are the same as in Fig. 4(a). By modulating all three parameters, $[x(t),y(t),J_{13}(t)]$, an alternative combination of state flips can be achieved. Indeed, as illustrated in Fig. 5(b), modulating $J_{13}$ enables the coupling of different pairs of the Riemann manifolds at time $t=T/2$, compared to the configuration shown in Fig. 4(b). Consequently, the following permutation $p=(1,2)(3,6)(4,5)(7,8)$ is implemented, as further confirmed by the calculated state fidelity in Fig. 5(d). We note that the same result can be also achieved when modulating the coupling $J_{23}(t)$ instead.
IV.3 Implementing a nontrivial permutation group in a three-qubit system
By choosing various combinations of the modulated qubit couplings $J_{mn}$, along with modulated parameters $(x,y)$ in Eq. (IV.1), six different permutations can be realized in the given three-qubit system. These permutations can be compactly expressed using cyclic notation as follows:
$p_{1}$ $=$ $(1,5)(2,6)(3,4)(7,8),$
$p_{2}$ $=$ (1,2)(3,6)(4,5)(7,8), if (J13) or $(1,2)(3,6)(4,5)(7,8),\text{ if $(J_{13})$ or $(J_{23})$},$
$p_{3}$ $=$ (1,5)(2,6)(3,7)(4,8), if $(1,5)(2,6)(3,7)(4,8),\text{ if $(J_{13},J_{23})$},$
$p_{4}$ $=$ (1,3)(2,6)(4,5)(7,8), if $(1,3)(2,6)(4,5)(7,8),\text{ if $(J_{12})$},$
$p_{5}$ $=$ (1,3)(2,4)(5,7)(6,8), if (J12,J13) or $(1,3)(2,4)(5,7)(6,8),\text{ if $(J_{12},J_{13})$ or $(J_{12},J_{23})$},$
$p_{6}$ $=$ (1,8)(2,6)(3,7)(4,5), if $(1,8)(2,6)(3,7)(4,5),\text{ if $(J_{12},J_{13},J_{23})$},$ (12)
where the parameters in the parenthesis $(J_{mn},\dots,J_{m^{\prime}n^{\prime}})$ indicate the qubit couplings that must be modulated to realize the corresponding permutation. We numerically confirm these findings in Appendix D.
The generators $p_{i}$ in Eq. (IV.3) form a small non-Abelian permutation group $\cal G$ with 576 elements, and whose group extension can be read as
${\cal G}=\Big{[}\left({\cal A}_{4}\times{\cal A}_{4}\right):{\cal C}_{2}\Big{]}:{\cal C}_{2}.$ (13)
The core of the group $\cal G$ is formed by the direct product of two alternating groups ${\cal A}_{4}$ of order 12, followed by nested semidirect products of the two cyclic groups ${\cal C}_{2}$ each of order 2.
The action of the group $\cal G$ on the eight elements, represented by the eight NHH eigenvectors, can be described as follows. The first direct product ${\cal A}_{4}\times{\cal A}_{4}$ represents the independent actions of two alternating groups ${\cal A}_{4}$, each acting on a disjoint set of four elements out of eight. Next, the cyclic group ${\cal C}_{2}$ acts on ${\cal A}_{4}\times{\cal A}_{4}$ by swapping the two ${\cal A}_{4}$ subgroups, effectively interchanging the two sets of 4 elements. Finally, an outer cyclic group ${\cal C}_{2}$ acts on the entire structure, providing an additional global transformation, such as inverting or permuting elements across the entire set.
Geometrically, the elements of the group $\cal G$ can describe the symmetries of chiral three-dimensional polytopes [54]. These geometric structures can naturally arise in the form of vertex-figures of higher-dimensional polytopes [54, 55].
Clearly, by using the generators $p_{i}$ in Eq. (IV.3), one can produce various permutations (576 in total) between the eigenvectors $\psi_{i}$. For instance, one can readily generate the following group elements of $\cal G$ written as
$p_{1}p_{3}p_{5}p_{1}p_{5}$ $=$ $(1,2)(3,4)(5,6)(7,8),$ (14)
$p_{1}p_{5}p_{1}$ $=$ $(1,8)(2,7)(3,6)(4,5).$ (15)
The group element in Eq. (14) represents a cyclic shuffling of eigenstates belonging to the two nearest energy manifolds, while the element in Eq. (15) corresponds to pairwise permutations between eigenstates whose energy levels are ‘mirror reflected’ to each other.
As such, a non-Abelian permutation group $\cal G$, acting on the eigenspace of the NHH, can be constructed through appropriate time modulations of the system parameters in Eq. (IV.1). This construction utilizes the nontrivial energy topology associated with the EPs of the pseudo-Hermitian Hamiltonian $H$, which effectively suppresses the occurrence of non-adiabatic transitions in the system’s dynamics.
V Discussion
It is evident that by considering larger systems with $(2n+1)>3$ qubits, more complex permutation groups can be implemented in the system dynamics. Importantly, the ability to realize permutation groups in the time-dependent NHH in Eq. (1) enables dynamical state transfer or swapping for any arbitrary state, since any state can be expressed as a superposition of the NHH eigenstates. This is in striking contrast with previous studies on dynamical winding around or in the vicinity of EPs, where non-adiabatic transitions prevent bijective mappings over NHH eigenstates due to chiral state switching [21, 1].
It is important to emphasize that when the initial state in the protocol described above is a superposition of the NHH eigenstates, the permuted eigenstates in that superposition accumulate an additional phase during state evolution. This phase generally comprises two components: the topological phase, arising from the encirclement of the EPs, and the dynamic phase. Such nontrivial phase accumulation can be further utilized in designing phase-dependent quantum protocols, akin to those employed in holonomic computation schemes [47].
The NHH of the form in Eq. (1) with effective complex magnetic fields can be experimentally realized, e.g., through a Naimark dilation method, where the given NHH constitutes a part of a larger Hermitian system [50, 56, 57, 58, 59]. Moreover, because the pseudo-Hermiticity of the NHH $\hat{H}$ (Eq. (1)) does not undergo spontaneous symmetry breaking, the exact form of the dilated Hermitian Hamiltonian can be readily found [58] (see also Appendix B for details). This method can be experimentally implemented using platforms based on nitrogen-vacancy centers in diamond [57, 59] or superconducting quantum processors [60, 61].
Alternatively, the NHH can be simulated in open quantum optical systems with hard-core bosons, where a fermionic system is effectively mapped to a bosonic one [62, 63]. In that case, the complex magnetic field can act as a complex driving coherent field, which can be emulated using postselection techniques based on homodyne detection [51]. However, in contrast to the ‘deterministic’ dilation method, that approach is inherently probabilistic, since it relies on a post-selection of trajectories that requires continuous monitoring of the environment. Furthermore, the experimental simulation of the NHH can also be performed in classical and quantum photonic setups consisting of a network of coupled waveguides, whose coupling topology is identified with the Fock space of the NHH [1].
The described protocol is robust against perturbations in the system that are sufficiently small to ensure that (i) the energy gap opening at the branch cut remains small enough that the state transfer is still feasible via Landau–Zener–Stückelberg–Majorana transitions [64], and (ii) the induced disorder does not trigger non-adiabatic transitions in the system dynamics. Indeed, in general, arbitrary disorder breaks the pseudo-Hermiticity of the NHH in Eq. (1), leading to a complex-valued spectrum and to system instabilities associated with this. We elaborate further on this in Appendix E.
Furthermore, variations in the orbiting trajectory of the magnetic field in the $(x,y)$ plane do not affect the resulting state permutations, as long as the dynamical loop encloses the EPs. The protocol also remains stable when the modulated interaction terms $J_{mn}\approx 0$ (but not exactly zero) at $t=T/2$ in Eq. (IV.1), meaning that the different Riemann sheet pairs do not precisely intersect at that moment. This robustness is again attributed to the Landau–Zener–Stückelberg–Majorana transitions between the corresponding energy manifolds, ensuring that the observed state permutations are still achievable [42].
We note that although in this work we specifically focused on the NHH in the form given in Eq. (1), other types of NHHs exhibiting various dynamic permutation groups in their eigenspace can also be identified in a similar manner. Our primary aim here was to highlight and describe the very possibility of the existence of such Hamiltonians that enable controlled and stable symmetric state switching in multiqubit systems via dynamical encirclement of EPs. Nevertheless, the formalism presented here can also be further extended to multiqubit systems with an even number of qubits. This could open up new possibilities for implementing genuine adiabatic quantum computing in non-Hermitian setups, e.g., realizing two-qubit quantum protocols in addition to the one- and three-qubit gates presented here.
VI Conclusion
In this work, we have demonstrated that various quantum state permutations can be realized in multiqubit non-Hermitian systems by dynamically encircling EPs. This is accomplished by specifically selecting a subspace of the system parameters where the NHH exhibits a real-valued spectrum. Dynamically winding around EPs within this subspace suppresses non-adiabatic transitions in the system dynamics, thereby enabling precise and controlled state permutations.
As an illustrative example, we studied an NHH describing an interacting three-qubit system. We showed that appropriately time-modulating the system parameter space allows implementing a nontrivial permutation group acting on the eigenstates of the given NHH. Furthermore, the group generators enable symmetric switching between any two eigenstates of the NHH, regardless of the winding direction.
The NHHs analyzed can be experimentally realized across various platforms, including solid-state systems such as superconducting qubits or nitrogen-vacancy centers in diamond using the dilation method. They can also be simulated in open quantum bosonic systems subject to quadrature fields postselection, or emulated in classical photonic setups through networks of coupled cavities or waveguides.
The ability to perform controlled state permutations, while dynamically encircling EPs, introduces new opportunities for quantum state generation and transfer. In particular, the method described here may lead to novel experimental approaches to performing holonomic quantum operations that are based on non-Hermitian quantum physics. Successful realizations of such schemes are likely to require additional corrective controls to manage spurious transitions when performing operations on quasi-diabatic timescales. The current findings thus highlight the potential of non-Hermitian systems to advance quantum information and communication protocols, opening up novel avenues for state control and manipulation.
Acknowledgements.
I.A. acknowledges support from Air Force Office of Scientific Research (AFOSR) Award No. FA8655-24-1-7376, from the Ministry of Education, Youth and Sports of the Czech Republic Grant OP JAC No. CZ.02.01.01/00/22_008/0004596, and from Grant Agency of the Czech Republic (Project No. 25-15775S). F.N. is supported in part by: the Japan Science and Technology Agency (JST) [via the CREST Quantum Frontiers program Grant No. JPMJCR24I2, the Quantum Leap Flagship Program (Q-LEAP), and the Moonshot R&D Grant Number JPMJMS2061], and the Office of Naval Research (ONR) Global (via Grant No. N62909-23-1-2074). S.K.O., P.L. and K.B.W. acknowledge support of AFOSR Multidisciplinary University Research Initiative (MURI) Award on Programmable systems with non-Hermitian quantum dynamics (Award No. FA9550-21-1-0202).
Appendix A Pseudo-Hermitian symmetry of the non-Hermitian Hamiltonian in Eq. (1)
Here we provide further details on pseudo-Hermiticity of the NHH in Eq. (1) in the main text. Namely, we want to show that one can find such a Hermitian operator $\eta$ for which one always has
$\eta\hat{H}\eta^{-1}=\hat{H}^{\dagger}.$ (16)
The operator $\eta$ can be chosen as
$\eta=\bigotimes\limits_{k}^{N}\xi,\quad\xi=\sum\limits_{i=1}^{2}|\nu_{i}\rangle\langle\nu_{i}|.$ (17)
The states $|\nu_{i}\rangle$ are the left eigenvectors of the single qubit Hamiltonian, i.e., $H_{q}^{\dagger}|\nu_{i}\rangle\equiv|\nu_{i}\rangle$, with $H_{q}=B_{x}\sigma_{x}+B_{y}\sigma_{y}$, and where $B_{x,y}$ are given in Eq. (1). That is, the operator $\xi$ is a metric operator of the qubit Hamiltonian $H_{q}$. We recall that the left eigenvectors $\nu_{1,2}$ are complex conjugates of the right eigenvectors in Eq. (7).
Because of this, the non-interacting part ($J=0$) of the NHH in Eq. (1), namely, $H_{\rm non}=\sum\limits_{k}^{N}\vec{{\boldsymbol{\sigma}}}^{k}\vec{\mathbb{B}}_{||}^{k}$, automatically fulfills the condition in Eq. (16), since the operator $\eta$ is a metric operator of this non-interacting term.
On the other hand, the interacting term $H_{\rm int}=\sum J_{mn}\hat{\sigma}_{z}^{m}\hat{\sigma}_{z}^{n}$ is invariant under the action of the operator $\eta$ in Eq. (17), i.e., $\eta\hat{H}_{\rm int}\eta^{-1}=H_{\rm int}$, which can be checked by straightforward calculations.
As such, one attains
$\eta\hat{H}\eta^{-1}=\eta\hat{H}_{\rm non}\eta^{-1}+\eta\hat{H}_{\rm int}\eta^{-1}=\hat{H}_{\rm non}^{\dagger}+\hat{H}_{\rm int}.$ (18)
Because the $\hat{H}_{\rm int}$ is Hermitian by definition, one arrives at the pseudo-Hermitian condition in Eq. (16).
Appendix B Realizing the non-Hermitian Hamiltonian in Eq. (1) via the dilation method
In this section, we review the key principles of the dilation method [50], which can be utilized for the experimental realization of the NHH in Eq. (1).
The core idea of the dilation method is to embed the Hilbert space of a given NHH into a larger space, where it becomes part of an extended Hermitian Hamiltonian. Consider the Schrödinger equation for a system governed by a NHH $H$ acting on an $n$-dimensional vector $|\psi\rangle$
$i\partial_{t}|\psi(t)\rangle=H|\psi(t)\rangle.$ (19)
To achieve that embedding one can define the following extended $2n$-dimensional state vector
$|\Psi\rangle=\begin{pmatrix}|\psi\rangle\\ D|\psi\rangle\end{pmatrix},$ (20)
where $D$ is a certain matrix to be defined later. The dynamics of this dilated vector $|\Psi\rangle$ is determined by a Hermitian Hamiltonian $\mathbf{H}$
$i\partial_{t}|\Psi(t)\rangle={\mathbf{H}}|\psi(t)\rangle,\quad{\mathbf{H}}={\mathbf{H}}^{\dagger}.$ (21)
By presenting this Hermitian Hamiltonian in the block matrix form ${\mathbf{H}}=\begin{pmatrix}h_{1}&h_{2}\\ h_{2}^{\dagger}&{h}_{3}\end{pmatrix}$, and combining Eqs. (19)–(21), one arrives to a system of differential equations
$h_{1}+h_{2}D$ $=$ $H,$
$h_{2}^{\dagger}+h_{3}D$ $=$ $i\dot{D}+DH.$ (22)
From the system of equations in Eq. (B) one attains a formal solution for the matrix $h_{1}$
$h_{1}=H+i\dot{D}^{\dagger}D-H^{\dagger}D^{\dagger}D+D^{\dagger}h_{3}D.$ (23)
By exploiting the property $h_{1}=h_{1}^{\dagger}$, one arrives to the equation
$i\partial_{t}\eta=H^{\dagger}\eta-\eta H,$ (24)
where we denoted $\eta=I+D^{\dagger}D$.
The equation Eq. (24) is precisely the equation for the metric of the NHH $H$ [65]. Indeed, by construction the $\eta$ is positive definite, and can be explicitly selected as given in Eq. (17). Because of the pseudo-Hermiticity of the NHH $H$, the metric $\eta$ in Eq. (24) is always ‘constant’ when written in the instantaneous left eigenbasis of the NHH. Moreover, with the choice in Eq. (17), the operator $(\eta-I)$ is non-negative, which immediately allows to find the ancillary matrix $D$ as
$D=U\sqrt{\eta-I},$ (25)
where $U$ is an arbitrary unitary matrix.
We stress that even for a time-dependent NHH, the metric retains the form given in Eq. (17), owing to the pseudo-Hermitian symmetry of the NHH in the whole system parameter space $(x,y,J)$. As such, one can readily construct a dilated Hermitian Hamiltonian whose subspace is constituted by the NHH from the knowledge of the metric $\eta$ in Eq. (17).
Appendix C Resolving the problem of exceptional points at infinity for single qubits in Eq. (3)
Here, we briefly address the apparent issue of spectral divergence at EPs in the non-interacting qubits NHH, according to Eq. (8). This problem can be resolved using the framework of complex projective spaces, which provides an elegant method to handle infinities. In particular, what appears as an infinity in a given space may correspond to a well-defined point in a higher-dimensional space.
To put it more formally, the procedure begins with identifying a higher-dimensional NHH that can be associated with a Grassmann manifold, which generalizes the concept of a projective space. Then, a particular class of this Grassmannian would corresponds to the lower-dimensional NHH, specifically the single-qubit NHH discussed in Sec. III.1.
Such a higher-dimensional $3\times 3$ NHH can be introduced as follows
$H^{\prime}=\begin{pmatrix}\gamma\cot\phi&\gamma&0\\ \gamma&-\gamma\cot\phi&0\\ 0&0&\dfrac{\sin\left({\rm Im}[\phi]\right)}{\sin\phi}\end{pmatrix},$ (26)
where $\gamma=\Delta\sin\left({\rm Re}[\phi]\right)$. The first two diagonal elements of the NHH are continuous functions on the plane $(x,y)$. The third element is also a continuous function except the interval $[x=0,|y|\leq 1]$, where it becomes rather a piecewise function, i.e., it is not differentiable but remains finite on that interval.
The NHH possesses an EP at the same point $(x=0,y=1)$ as the NHH of a single qubit considered in Sec. III.1. Moreover, the spectrum of the NHH $H^{\prime}$ in Eq. (26) is finite in any given finite region $(x,y)$ including the EP. These eigenvalues read
$\lambda_{1,2}=\pm\gamma(\sin\phi)^{-1},\quad\lambda_{3}=\sin\left({\rm Im}[\phi]\right)(\sin\phi)^{-1}.$ (27)
The second-order EP is characterized by the coalescence of two eigenvectors associated with the degenerate eigenvalues
$\lambda_{1}^{\rm EP}=\lambda_{2}^{\rm EP}=0.$ (28)
Notably, while the third eigenvalue is also $\lambda_{3}=0$ at the EP, its corresponding eigenvector remains independent of the other two eigenvectors.
Figure 6: This Figure is an extension of Figs. 3(d) and 4(d), i.e., the calculated state transfer fidelity $|\langle\psi_{k}|\psi(t)\rangle|^{2}$ between evolving state $|\psi(t)\rangle$ and one of the eight initial eigenstates $|\psi_{k}\rangle$ of the three-qubit NHH. The panels in the plot are related to different combinations of the modulated system parameters $J_{mn}$ in Eq. (IV.1), and which thus correspond to permutation generators in Eq. (IV.3). The panels (a),(b),(c) and (d) correspond to the permutation generators $p_{3}$, $p_{4}$, $p_{5}$ and $p_{6}$, respectively. The remaining system parameters are the same as in Figs. 3(d) and 4(d).
Now, we can associate subspaces of this $3\times 3$ NHH with a Grassmanian ${G}_{1,3}$, which is the projective space ${\mathbb{C}P}^{2}$ [66]. In the case under consideration, we define the projective space ${\mathbb{C}P}^{2}$ with the coset $cH^{\prime}$, where $c\in{\mathbb{C}}-\{0\}$.
One of the charts of the this projective space can be defined as $\dfrac{\sin\phi}{\sin\left({\rm Im}[\phi]\right)}H^{\prime}$, obtained by normalizing $H^{\prime}$ by its last diagonal element. The representative of this chart with inhomogeneous coordinates can then be expressed via the $2\times 2$ matrix $H=\alpha\begin{pmatrix}\cos\phi&\sin\phi\\ \sin\phi&-\cos\phi\end{pmatrix}$, which corresponds to the NHH of the single qubit discussed in Sec. III.1.
In other words, through this projection, the well-defined EP in Eq. (28) is sent to infinity in Eq. (8). The latter corresponds to the situation when the projection of the EP in ${\mathbb{C}}^{3}$, determined by the $3\times 3$ matrix $H^{\prime}$ to the plane, defined by the $2\times 2$ matrix $H$, is parallel to this plane.
Thus, the projective space associated with the NHH $H$ provides a framework to effectively map the EP in Eq. (28) to infinity, according to Eq. (8). We reserve the discussion of experimental implementation of this mathematical approach for future studies.
Appendix D Calculated fidelity for other state permutations according to Eq. (IV.3)
In Figs. 3(d) and 4(d) we plotted the fidelity $|\langle\psi_{k}|\psi(t)\rangle|^{2}$ between evolving state $|\psi(t)\rangle$ and one of the eight initial eigenstates $|\psi_{k}\rangle$ of the three-qubit NHH that correspond to the first generators of the permutation group $\cal G$ in Eq. (IV.3). For completeness, here in Fig. 6 we plot the fidelity of the remaining permutation generators, thus confirming their form in Eq. (IV.3).
We also summarize various eigenstate flips produced by the permutation generators $p_{i}$ in Table 1. From this Table it is clear that one can always find such a combination of modulated system parameters in Eq. (IV.1) that enables a state transfer between any two given eigenstates in one or more full cycles $T$.
Table 1: Table of the state permutations according to Eq. (IV.3). Starting from an initial system state $\psi_{k}$ (listed in the first column), one can transition to any target eigenstate $\psi_{j}$ (listed in the first row) by selecting an appropriate trajectory (expressed via a permutation generator $p_{i}$ in Eq. (IV.3)) that winds or spirals around the exceptional point.
$\psi_{\rm in}$ $\psi_{\rm out}$ $\psi_{1}$ $\psi_{2}$ $\psi_{3}$ $\psi_{4}$ $\psi_{5}$ $\psi_{6}$ $\psi_{7}$ $\psi_{8}$
$\psi_{1}$ $p_{2}$ $p_{4},p_{5}$ $p_{1},p_{3}$ $p_{6}$
$\psi_{2}$ $p_{2}$ $p_{5}$ $p_{1},p_{3},p_{4},p_{6}$
$\psi_{3}$ $p_{4},p_{5}$ $p_{1}$ $p_{2}$ $p_{3},p_{6}$
$\psi_{4}$ $p_{5}$ $p_{1}$ $p_{2},p_{4}$ $p_{3}$
$\psi_{5}$ $p_{1},p_{3}$ $p_{2},p_{4}$ $p_{5}$
$\psi_{6}$ $p_{1},p_{3},p_{4},p_{6}$ $p_{2}$ $p_{5}$
$\psi_{7}$ $p_{3},p_{6}$ $p_{5}$ $p_{1},p_{2},p_{4}$
$\psi_{8}$ $p_{6}$ $p_{3}$ $p_{5}$ $p_{1},p_{2},p_{4}$
Appendix E Effects of perturbations on dynamical state permutations in three-qubit system
Here, we briefly examine the impact of perturbations in the three-qubit NHH on the state permutations discussed in Sec. IV.3. Generally, arbitrary perturbations to the NHH in Eq. (1) lead to a complex spectrum with a non-zero imaginary part, which plays a critical role in triggering NATs in the system dynamics [21]. For sufficiently large perturbations, the permutation symmetry of the states is broken, resulting in chiral state dynamics. However, when perturbations remain small, the symmetry in state flips can be preserved.
Without loss of generality, we assume that a three-qubit NHH $\hat{H}$ in Eq. (1) is perturbed as follows
$H\to H+\epsilon\sigma_{z}^{1}.$ (29)
That is, there is a non-zero magnetic field perturbation applied to the first qubit along the $z$-axis. The effects of this perturbation on the state transfer protocol are shown in Fig. 7 and Fig. 8. Figure 7 (Fig. 8) corresponds to the unperturbed case shown in Fig. 3 (Fig. 4). As seen in both figures, the state evolution under perturbation starts to differ based on the winding direction, with this difference growing as the disorder increases, eventually leading to state evolution instabilities and NATs (not shown).
However, as demonstrated in Figs. 7 and 8, for small levels of disorder ($\epsilon\approx 0.01$), the desired state permutations remain feasible. By appropriately tuning the system parameters, the state transfer fidelity can still be kept at high rates.
Figure 7: State transfer fidelity for the perturbed three-qubit NHH, according to Eq. (1) and Eq. (29) for a perturbation $\epsilon=0.01$. This figure corresponds to the unperturbed case shown in Fig. 3(d). The upper (lower) row of panels demonstrates the scenario when the dynamical winding is performed in the counterclockwise $+\omega$ (clockwise $-\omega$) direction. The time period is chosen as $T=50$ [arb. units], the remaining system parameters are the same as in Fig. 3.
Figure 8: State transfer fidelity for the perturbed three-qubit NHH, according to Eq. (1) and Eq. (29) for a perturbation $\epsilon=0.01$. This figure corresponds to the unperturbed case shown in Fig. 4(d). The upper (lower) row of the panels shows the scenario when the dynamical winding is performed in the counterclockwise $+\omega$ (clockwise $-\omega$) direction. The time period is chosen as $T=120$ [arb. units], and $g_{0}=1.5$, the remaining system parameters are the same as in Fig. 4.